### v2

# Stellar Spectroscopy Using Diffraction Grating, CMOS Monochrome Sensor, and Reflecting Telescopes

Abstract
We present the design and testing of a compact, low-cost stellar spectrometer developed for undergraduate and outreach applications. The instrument employs a 600 lines/mm diffraction grating, a CMOS monochrome sensor, and a 3D-printed mount integrated with reflecting telescopes. Calibration was performed using helium emission sources in the laboratory and Vega as a spectrophotometric standard, supported by a custom Python-based image-processing pipeline for wavelength calibration and spectral stacking. The spectrometer successfully recorded usable spectra of bright stars including Vega, Sirius, Procyon, Capella, and Betelgeuse, covering spectral types A through M. The results demonstrate that meaningful stellar spectroscopy can be achieved with accessible, low-cost equipment, providing a practical framework for student-led astronomical instrumentation projects.
1 Introduction
Since the 19th century, stellar spectroscopy has revolutionized our understanding of astrophysics, beginning with Fraunhofer’s discovery of dark absorption lines in the solar spectrum (fraunhofer1814; Kirchoff1860FL; Kirchoff1860Sun). By the early 20th century, Annie Jump Cannon’s Harvard classification and Cecilia Payne’s work on stellar atmospheres established spectroscopy as the foundation of stellar astrophysics, linking spectral signatures to temperature, chemical composition, and stellar evolution (1925Payne; maury1897spectra).
Today, high-resolution spectroscopy remains a cornerstone of observational astrophysics, enabling precise measurements of stellar abundances, radial velocities, and the detection of exoplanetary atmospheres (2005DavidG; 2005Martin). Recent usage of HIRES on the Keck I Telescope, employs a cross-dispersed echelle-grating spectrograph, achieving spectral resolutions between 25000 to 85000 across a wavelength range of 0.3 to 1.0 microns (HIRES). This high resolution allows for precise measurements of stellar parameters and has been instrumental in various astrophysical studies. Inspired by these sophisticated instruments, our project aims to construct a compact, low-cost diffraction-grating spectrometer integrated with a reflecting telescope and a CMOS sensor.
While operating at a lower resolution, the instrument embodies the similar fundamental techniques, providing a hands-on platform to observe and analyze stellar spectra across multiple spectral classes. This approach not only offers practical experience in astronomical instrumentation but also contributes to the growing field of accessible astronomical spectroscopy. Using a 600 lines/mm diffraction grating with a CMOS monochrome sensor, and a custom 3D-printed mount coupled to an 11-inch Schmidt-Cassegrain telescope (Figure 1), we constructed a spectrometer capable of dispersing starlight into well-defined spectra. Our targets – Sirius (A1V, the brightest star in the night sky), Vega (A0V, a standard photometric reference), Procyon (F5IV, a nearby subgiant), Capella (G8III + G0III, a quadruple system with a primary-pair of giants), and Betelgeuse (M2Ia, a red supergiant) (Gray_2003_Sir_Vega; Perkins_BG_cap; procyon0; Torres_2015_Capella_bin) – were chosen to span a wide range of stellar types and evolutionary stages. Each of these stars carries historical significance in spectroscopy: Vega as the flux calibration standard, Sirius as an early benchmark in line-profile studies, and Betelgeuse as one of the first stars with a resolved stellar disk.
The overarching aim was to demonstrate that stellar spectra from such diverse stars could be recorded and calibrated while being feasible at the institute or undergraduate research level, broadening access to hands-on observational experience. In doing so, the project highlights both the potential and limitations of low-cost spectroscopy, while offering an accessible framework for expanding its application in both research and education.
2 Methods
2.1 Stellar Spectroscopy Principles
Every star carries its own spectral fingerprint - an intricate code of light that reveals its hidden physical and chemical identity. Through stellar spectroscopy, we decode this light to study the temperature, composition, and motion of astrophysical bodies far beyond our reach. When a star emits radiation, its photosphere acts as a continuous light source, while the cooler gases in its outer layers absorb specific wavelengths corresponding to electronic transitions of various atoms and ions. These absorbed wavelengths appear as dark absorption lines within the continuous spectrum, each marking the presence of an element such as hydrogen, calcium, or iron.
The systematic study of these spectral lines forms the foundation of stellar classification and atmospheric modeling. For instance, A-type stars like Vega and Sirius display striking hydrogen Balmer lines, indicative of high photospheric temperatures ($\sim$10,000 K). In contrast, G- and K-type stars such as Procyon or Capella exhibit stronger metallic lines, while M-type giants like Betelgeuse reveal molecular absorption bands from titanium oxide in the red and near-infrared regions. These differences can trace stellar temperature and ionization states to yield information about stellar evolution and surface gravity. However, terrestrial observations face challenges such as Doppler broadening, pressure broadening, and telluric absorption from Earth’s atmosphere - each requiring careful calibration to preserve the true stellar signature. By studying these features, even small-scale spectroscopic experiments can capture the essence of stellar physics that drives modern astrophysical research.
2.2 Diffraction Grating Theory
For this project, we use a diffraction grating, which is an optical grating that diffracts light into its constituent wavelengths into different directions. Unlike a prism, which relies on refraction and introduces chromatic aberrations, a grating uses the principle of interference between light waves diffracted by a regular array of slits or grooves. This allows for sharper, more evenly spaced spectral lines and enhanced wavelength resolution, an essential advantage for observing subtle stellar absorption features. The choice of a grating-based spectrometer thus provides both affordability and scientific accuracy, making it particularly well-suited for educational and research applications in stellar spectroscopy.
A transmission diffraction grating disperses incident light according to the grating equation:
$d\sin\theta=n\lambda$ (1)
where $d=1/N$ is the grating spacing (N = 600 lines/mm for our grating), $\theta$ is the diffraction angle, $n$ is the diffraction order, and $\lambda$ is the wavelength. For optimal spectral resolution with minimal overlapping orders, we operated in first-order diffraction ($n=1$), yielding angular dispersion of approximately $12.1\degree$ to $24.8\degree$ across the visible range (350-700 nm).
Figure 1: Left: Skyris 236M CMOS monochrome sensor used for spectral imaging. Right: 11-inch Celestron CPC Deluxe 1100 HD telescope with computerized tracking, used for primary stellar observations.
Figure 2: Left: Ideal geometric alignment showing the diffraction grating, sensor position, and angular geometry for first-order diffraction. The distance $l$ between grating and sensor was optimized to capture the full visible spectrum. Right: 3D model of the custom 3D-printed spectrometer mount designed in SolidWorks. The mount securely holds the diffraction grating and CMOS sensor while interfacing with standard telescope eyepieces.
2.3 Instrument Design and Construction
Our spectrometer consists of three main components: a 600 lines/mm transmission diffraction grating, a Skyris 236M CMOS monochrome sensor, and a custom 3D-printed mount that attaches to standard telescopes. The key specifications of our equipment are summarized in Table 1.
Component Specification
Diffraction grating 600 lines/mm transmission
CMOS sensor Skyris 236M monochrome
Sensor dimensions 5.44 mm $\times$ 3.67 mm
Primary telescope 11-inch CPC Deluxe 1100 HD
Secondary telescope 8-inch SkyWatcher Quattro 200P
Slit width range 1.0 - 1.5 mm
Wavelength coverage 350 - 700 nm
Table 1: Key equipment specifications used in our stellar spectrometer
The optimal sensor-grating geometry was derived from diffraction theory to capture the complete visible spectrum (350-700 nm) across the sensor width. We focused on first-order diffraction ($n=1$) because higher orders would require prohibitively large diffraction angles that exceed our compact sensor geometry, demanding a much larger sensor or impractically small grating-to-sensor distance, making it physically impossible. For the desired wavelength range $\lambda_{1}$ to $\lambda_{2}$, the corresponding diffraction angles are:
$\theta_{1}=\sin^{-1}(\lambda_{1}N)\quad\text{and}\quad\theta_{2}=\sin^{-1}(\lambda_{2}N)$ (2)
While the required distance $l$ between grating and sensor is given by:
$l=\dfrac{s}{2\sim\tan\left(\dfrac{\theta_{2}-\theta_{1}}{2}\right)}$ (3)
where $s$ is the sensor width (5.44 mm). With $\theta_{1}=12.1\degree$ and $\theta_{2}=24.8\degree$, we calculated $l\approx 25$ mm, positioning the spectrum centrally on the sensor.
The spectrometer housing was designed in SolidWorks and fabricated via fused deposition modeling using PLA plastic. The mount integrates a standard 1.25 inch telescope eyepiece adapter, an adjustable grating holder for fine angular tuning, and a rigid sensor fixture maintaining a 15 mm focal distance from the camera interface. This configuration ensures precise optical alignment and ease of integration with both the 8-inch and 11-inch telescopes used for observations. Figure 2 illustrates the 3D model and the ideal optical configuration.
2.4 Slit System
During preliminary tests using a collimated helium lamp placed at large distances (hundreds of meters), we observed that spectral lines appeared broadened when viewed through the telescope. Although both stars and the lamp act as point sources at infinity, telescope optics and atmospheric seeing cause them to form finite disks. Consequently, the stellar image was spread over multiple sensor rows, producing broadened spectral lines and reducing effective resolution. To prevent this, we used a narrow entrance slit to confine the incoming light and preserve fine spectral detail.
We implemented an adjustable precision-blade slit system with widths between 1.0 and 1.5 mm, positioned ahead of the spectrometer grating. The slit acts as a spatial filter, admitting light from a narrow region of the stellar image into the spectrometer. Selecting the optimal slit width required balancing three competing factors: spectral resolution, light throughput, and diffraction effects.
The 1.0-1.5 mm range provided the best compromise for our sensor dimensions (5.44 mm width) and telescope focal ratios (f/4 to f/10), maintaining adequate light levels for reasonable exposure times while preserving the resolution needed to resolve individual absorption features.
2.5 Calibration and Testing
The spectrometer was first validated under controlled laboratory conditions using helium emission lamps as wavelength standards. The helium lamp provided six prominent lines (447.1–706.5 nm) across the visible range. Initial alignment with a pinhole collimator produced internal reflections, which were mitigated by optical baffling and path optimization. Pixel positions of the identified emission peaks were matched to known wavelengths to obtain a polynomial wavelength calibration:
$\lambda(\text{nm})=-0.000079x^{2}-0.0381x+711.502,$ (4)
where $x$ denotes the pixel column. Although the grating equation (Equation 1) implies a sinusoidal relation, a polynomial fit was preferred due to minor optical misalignments and the near-linear pixel-to-angle projection across the small sensor field (Figure 3).
Figure 3: Wavelength calibration using helium emission lines. The quadratic fit relates pixel position to wavelength, with residuals <<1 nm across the visible range.
Following laboratory calibration, we tested the spectrometer on two reflecting telescopes: an 8-inch (f/4) and an 11-inch (f/10). The larger aperture increased light-gathering power and allowed shorter exposures, but results in broadened stellar images at the focal plane, reducing resolution. We therefore used an adjustable slit as detailed earlier, and computerized tracking on the 11-inch to focus on targets for sequential exposures and stacking.
Using helium lamp lines in both the laboratory and at the telescope, we also measured the instrumental resolving power. The fitted FWHM is nearly constant in detector pixels (mean $\approx 101\pm 8$ px), while the dispersion is strongly wavelength dependent—ranging from $\approx 0.051$ nm/px ($0.51$ Å/px) at 706.5 nm to $\approx 0.292$ nm/px ($2.92$ Å/px) at 447.1 nm—so the resolving power declines from $R\approx 142$ (706.5 nm) to $R\approx 16$ (447.1 nm). This pixel-limited broadening, together with the wavelength-dependent dispersion, means narrow lines are more blended in the blue, whereas continuum slopes and broad molecular bands remain well preserved; accordingly, we bin the spectra to improve SNR for the objectives of this work.
2.6 Data Processing Pipeline
We developed a comprehensive Python-based analysis pipeline to systematically process raw spectral images and extract calibrated stellar spectra. The pipeline integrates image preprocessing, wavelength calibration, spectral alignment, and stacking procedures.
Image Preprocessing
Raw FITS images from the CMOS sensor were preprocessed to isolate the spectral signal from detector noise. We summed pixel intensities along each row (perpendicular to dispersion direction) and fitted a Gaussian profile to identify the spectral trace center and width. The central region was extracted using a window defined by the full-width-at-half-maximum (FWHM) of the Gaussian fit of the vertical span, effectively rejecting the rest of the image contaminated by other sources (eg, hot pixels, light pollution, etc.) outside the primary spectral trace.
Wavelength Calibration Using Vega
Vega ($\alpha$ Lyrae, spectral type A0V) served as our primary spectrophotometric standard. We adopted the Hubble Space Telescope absolute spectrophotometry of Vega from bohlin2014 as our reference spectrum. This choice addresses the lack of manufacturer-provided sensor response curves by implicitly calibrating instrumental sensitivity through comparison with a well-characterized stellar standard.
Prominent hydrogen Balmer absorption lines (H-$\alpha$ at 6563 Å, H-$\beta$ at 4861 Å) were identified in raw Vega observations and used as wavelength fiducials. To construct a wavelength-dependent sensitivity function $w(\lambda)$, we related observed intensities $I_{\text{obs}}$ to model intensities $I_{\text{model}}$ via:
$w(\lambda)=\frac{I_{\text{model}}(\lambda)}{I_{\text{obs}}(\lambda)}$ (5)
The sensitivity function was smoothed using Savitzky-Golay filtering and interpolated across the full wavelength range to enable calibration of subsequent stellar observations. Figure 4 shows the recorded Vega spectrum after wavelength calibration.
Figure 4: Recorded Vega spectrum with the blue error bands representing standard deviations from stacked images, demonstrating the robustness of our data reduction pipeline.
Spectral Alignment and Image Stacking
Atmospheric seeing, tracking errors, and instrumental flexure introduce sub-pixel shifts between sequential exposures. To optimally align spectra before stacking, we employed $\chi^{2}$ minimization across identified spectral markers (typically H-$\alpha$ and H-$\beta$ absorption lines):
$\chi^{2}=\sum_{i=1}^{N}\sum_{j=1}^{N^{\prime}}\frac{(\lambda_{i,j,\text{obs}}-\lambda_{j,\text{expected}})^{2}}{\lambda_{j,\text{expected}}}$ (6)
where $N$ is the number of images, $N^{\prime}$ is the number of spectral markers, and shift parameters are optimized for each image to minimize total $\chi^{2}$.
After alignment, spectra were binned into 1 Å wavelength intervals. For each bin, we calculated the weighted mean intensity and standard deviation across all images, yielding a high signal-to-noise final spectrum with associated uncertainty estimates - a feature uncommon in amateur spectroscopy but essential for quantitative analysis. Figure 5 illustrates the spectral marker identification process used for alignment.
Figure 5: Identification of H-$\alpha$ and H-$\beta$ spectral markers in Vega observations. The algorithm identifies absorption line minima across multiple frames, enabling precise spectral alignment via $\chi^{2}$ minimization before stacking.
3 Results
3.1 Spectral Features by Stellar Type
Star Spectral Class Key Features
Vega A0V Strong H lines
Sirius A1V Strong H lines
Procyon F5IV Metallic lines emerging
Capella G8III + G0III Strong metallic lines
Betelgeuse M2Ia TiO bands, Na doublet
Table 2: Stellar targets observed with our spectrometer, covering spectral classes A through M
Our spectrometer successfully captured calibrated spectra from five bright stars spanning spectral classes A through M (Table 2). Typically 100 individual exposures per target were processed through the pipeline and stacked to produce final spectra with wavelength coverage from approximately 3500 to 7500 Å.
(a) Vega (A0V)
(b) Sirius (A1V)
Figure 6: A-type stellar spectra. (a) Vega (A0V) spectrum showing strong hydrogen absorption lines characteristic of A-type stars. (b) Sirius (A1V) spectrum with similar hydrogen features but slightly enhanced metallic lines. Both spectra display characteristic A-type stellar properties with prominent Balmer series absorption.
(a) Procyon (F5IV)
(b) Capella (G8III + G0III)
(c) Betelgeuse (M2Ia)
Figure 7: Stellar spectra showing the progression from F to M spectral types. (a) Procyon (F5IV) spectrum showing weakened hydrogen lines and emerging metallic absorption features. (b) Capella (G8III + G0III) spectrum dominated by metallic absorption lines, particularly strong Ca II H and K lines. (c) Betelgeuse (M2Ia) spectrum showing minimal hydrogen absorption and strong molecular features, with prominent Na I doublet and broad TiO absorption bands characteristic of cool M-type supergiant atmospheres.
A-Type Stars (Vega and Sirius)
Vega and Sirius (Figure 6) exhibited the expected strong hydrogen Balmer absorption lines characteristic of A-type stellar atmospheres. We identified H-$\gamma$ (4340 Å), H-$\beta$ (4861 Å), and H-$\alpha$ (6563 Å) with line depths of 15–25% of the continuum (Figure 6). Our Vega spectrum showed excellent agreement with HST reference data, validating our calibration approach. Sirius displayed slightly stronger metallic features (Na I doublet at 5890/5896 Å), consistent with its A1V classification versus Vega’s A0V type. Telluric absorption bands (O2 at 6867–6884 Å, H2O at 7167–7350 Å) were visible in both spectra, as expected for ground-based observations.
F-Type Star (Procyon)
Procyon (F5IV) (Figure 7) showed weakening hydrogen lines relative to A-type stars, with H-$\beta$ depth reduced to $\sim$10%. Metallic absorption lines became more prominent, including Ca-II, H and K lines (3968/3933 Å), Fe-I lines (multiple features at 4450–6200 Å), and enhanced Na I doublet (Figure 7). This trend toward metal-line dominance at lower temperatures aligns with Harvard classification criteria.
G-Type Star (Capella)
Capella (G8III + G0III) (Figure 7) exhibited further weakening of hydrogen features and strengthening of metallic lines. The Ca-II, H and K resonance lines appeared particularly strong, as expected for evolved giant stars with extended atmospheres. Multiple Fe-I and Fe-II lines dominated the blue-green spectral region (4000-5500 Å, Figure 7). The overall spectral energy distribution peaked in the yellow-green region, consistent with Capella’s composite surface temperature of $\sim$4970 K(Torres_2015_Capella_bin).
M-Type Star (Betelgeuse)
Betelgeuse (M2Ia) (Figure 7) displayed the coolest spectrum with barely detectable hydrogen absorption and strong molecular absorption features. The Na-I doublet appeared prominently in emission or partial absorption-filling, characteristic of extended supergiant atmospheres (Figure 7). While our spectral resolution was insufficient to fully resolve TiO molecular band structure, we observed broad absorption features centered near 5280 Å consistent with the $\gamma$ band of TiO. The red-peaked spectral energy distribution confirmed Betelgeuse’s classification as a red supergiant.
3.2 Comparison with Observational Data
Direct comparison of our Sirius spectrum with HST observations revealed remarkable correspondence in major absorption features despite the difference in aperture and detector sophistication (Figure 8). Line positions agreed within our 1 Å binning resolution, and relative line strengths matched to within 20-30%, validating both instrumental calibration and data reduction procedures. The primary differences were decreased signal-to-noise in our data and stronger telluric contamination, both expected for small-aperture ground-based observations.
(a) Vega comparison
(b) Sirius comparison
Figure 8: Comparison of our calibrated spectra (red) with HST observations and model (blue). (a) Vega spectrum comparison showing excellent agreement in major absorption features between our instrument and HST model. (b) Sirius spectrum comparison demonstrating similar correspondence with HST data. Major absorption features align well, validating our calibration and data reduction methods.
4 Discussion
Our low-cost stellar spectrometer successfully demonstrated the feasibility of spectroscopy using accessible equipment. The instrument achieved wavelength calibration accuracy better than 1 nm across the visible spectrum and captured diagnostically useful spectra spanning spectral classes A through M. The excellent agreement between our Vega and Sirius observations with HST reference data validates both our calibration methodology and data reduction pipeline.
The choice of Vega as a spectrophotometric standard proved particularly effective, addressing the absence of manufacturer-provided sensor response curves through implicit calibration. This approach enabled quantitative comparison with established stellar standards while maintaining accessibility for educational applications. The $\chi^{2}$ minimization technique for spectral alignment, combined with 1 Å binning and uncertainty propagation, produced results remarkably similar to HST observations, which is an excellent result for amateur spectroscopy and demonstrates the potential of low-cost instrumentation.
Several design choices proved critical to instrument performance. The 600 lines/mm grating provided optimal dispersion for our small sensor, while first-order diffraction avoided spectral overlap complications. The precision slit system (1.0-1.5 mm) effectively balanced light throughput against spectral resolution, though narrower slits introduced problematic Fresnel diffraction artifacts. The 3D-printed mount’s mechanical stability and telescope interface design enabled consistent results across multiple observing sessions.
The measured resolving power reveals important instrument characteristics. The FWHM remains approximately constant in pixel units (mean $\approx 101\pm 8$ px), indicating that spectral broadening arises primarily from the slit/optical image size rather than wavelength-dependent effects. However, wavelength-dependent dispersion, ranging from $\approx 0.051$ nm/px ($0.51$ Å/px) at 706.5 nm to $\approx 0.292$ nm/px ($2.92$ Å/px) at 447.1 nm, causes resolving power to decline from $R\approx 142$ in the far red to $R\approx 16$ in the blue. Narrow lines are thus more blended in the blue, though broad continuum trends and molecular bands (evolving on scales of tens to hundreds of Å) remain well preserved across 3500–7500 Å. For educational use, we recommend reporting wavelength-dependent $R(\lambda)$ values, presenting raw and binned data to illustrate the S/N versus resolution trade-off, and emphasizing the pixel-limited nature of the instrument. These considerations set realistic expectations: continuum fitting, synthetic color indices, and coarse classification remain appropriate, whereas line-profile analysis or radial-velocity work lie beyond the instrument’s capabilities.
Key limitations include restricted wavelength coverage (3500-7500 Å), preventing near-UV and near-IR observations, and insufficient resolution for fine spectral structure. Telluric contamination complicated analysis near atmospheric absorption lines. The magnitude limit (V << 3) restricts targets to bright stars, though this aligns well with educational objectives.
We have successfully constructed and characterized a compact, low-cost stellar spectrometer that produced reliable results across multiple spectral classes. Observations of five stars spanning spectral types A0V through M2Ia revealed the expected progression of spectral features: strong hydrogen absorption in A-type stars, the emergence of metallic lines in F-type stars, prominent Ca-II, H and K features in G-type giants, and molecular absorption bands in M-type supergiants. The image-processing pipeline, which incorporated $\chi^{2}$-based spectral alignment and uncertainty estimation, yielded consistent and reproducible results. Overall, this work demonstrates the feasibility of conducting meaningful stellar spectroscopy with accessible equipment, emphasizing its potential as a practical tool for educational and undergraduate research in observational astronomy.
Acknowledgements
We thank Dr. Kartik Senapati for his supervision, guidance, and encouragement throughout this project, and Dr. Gunda Santosh Babu for valuable feedback on observational procedures. We acknowledge the technical support from Mr. Rudranarayan Mohanty and Mr. Sakthivel V. A. for laboratory resources and equipment access. We also gratefully acknowledge the Integrated Lab (Physics) setting at School of Physical Sciences, NISER for providing the opportunity and resources to propose, develop, and undertake this undergraduate experimental project. This work would not have been possible without the generous support of the NISER Astronomy Club [^1] NISER Astronomy Club astroclub-niser.github.io (NAC), which provided access to the Skyris 236M CMOS sensor and both the 8-inch and 11-inch reflecting telescopes. We are particularly grateful to Vasanth Kashyap for assistance with spectrometer design and 3D modeling, and Abha Vishwakarma for insightful discussions on data analysis methods. The NISER RoboTech Club [^2] NISER Robotech Club rtcniser.github.io manufactured the spectrometer mount via 3D printing. We thank the NAC members for assistance with telescope setup during nighttime observing sessions.
Data Availability
The raw spectral data and the processed spectral datasets obtained using the telescope for stellar observations and calibration sources are available from the authors upon reasonable request. The data processing pipeline and analysis code used in this work are publicly available in our GitHub repository at https://github.com/niti21/Stellar-Spectrometry.